write an equation for the nth term of the geometeric sequence 9/16, 3/8, 1/4.
Find the seventh term of this sequence.

Answer by unlockmath(1688)

Hello,
First find the ratio which is:
3/8 over 9/16 = 2/3
7th term=9/16(2/3)^7-1
7th term=4/81
nth term=9/16(2/3)^n-1
Make sense?
RJ
www.math-unlock.com